OpenStudy (anonymous):

24% of what number is 19.92?

OpenStudy (anonymous):

19.92=24(1/100)x 1992/24 = x 83=x

OpenStudy (anonymous):

24% of X is 19.92 since x is the unkown quantity thats we're looking for. So 24%x=19.92. Isolate x by dividing bothe sides by 24. So 24/24x=19.92/24 wich equals 0.83 then times .83 by 100